can you show step by step how to do this problem -13a+6-4c-11+5a-2c
Answer by Edwin McCravy(20086)


-13a+6-4c-11+5a-2c

Combine the red terms. They are like terms because they both have
a's:    -13a+5a = -8a.

Combine the purple terms. They are like terms because they both are just 
plain numbers with no letter:    +6-11 = -5.

Combine the green terms. They are like terms because they both have 
c's:    -4c-2c = -6c.
 
Final answer:  

-8a-5-6c.

Edwin